### v1

# APOLLO Blender: A Robotics Library for Visualization and Animation in Blender

Abstract
High-quality visualizations are an essential part of robotics research, enabling clear communication of results through figures, animations, and demonstration videos. While Blender is a powerful and freely available 3D graphics platform, its steep learning curve and lack of robotics-focused integrations make it difficult and time-consuming for researchers to use effectively. In this work, we introduce a lightweight software library that bridges this gap by providing simple scripting interfaces for common robotics visualization tasks. The library offers three primary capabilities: (1) importing robots and environments directly from standardized descriptions such as URDF; (2) Python-based scripting tools for keyframing robot states and visual attributes; and (3) convenient generation of primitive 3D shapes for schematic figures and animations. Together, these features allow robotics researchers to rapidly create publication-ready images, animations, and explanatory schematics without needing extensive Blender expertise. We demonstrate the library through a series of proof-of-concept examples and conclude with a discussion of current limitations and opportunities for future extensions.
I Introduction
Clear and compelling visuals are central to the communication of robotics research. Whether in the form of figures for papers, animations for submission videos, or explanatory graphics for presentations, the ability to convey complex concepts visually is critical for both technical audiences and broader outreach. While modern 3D graphics software such as Blender provides powerful tools for creating high-quality renderings and animations, these platforms are rarely designed with robotics in mind. As a result, researchers often face steep barriers when attempting to generate robotics-specific visuals, requiring significant effort to learn software workflows, manually configure robots, and avoid visual inconsistencies or bugs.
Existing solutions within the robotics community only partially address this challenge. Game engines such as Unity or Unreal have been adopted for real-time simulation and interactive demonstrations, but they require extensive development time to achieve high-quality renderings. Dedicated simulators such as Gazebo, CoppeliaSim, or Webots offer robust physics engines and standard robot description support, yet they typically prioritize accurate dynamics over visualization fidelity. Specialized visualization tools, meanwhile, are often restricted to educational use cases or narrow application domains, limiting their flexibility and visual quality. Across this spectrum, there remains a gap between tools optimized for physical accuracy and tools optimized for high-quality static and animated renderings.
In this work, we present APOLLO Blender, a lightweight software library that lowers this barrier by providing robotics-focused scripting tools within Blender. Our goal is not to replicate the full functionality of physics simulators or game engines, but rather to make it straightforward for researchers to create static images, animations, and schematic illustrations that are both visually appealing and robotics-aware. The library centers on three key features: importing robots and environments from standardized descriptions such as URDF, keyframing robot configurations and appearance attributes, and generating common 3D primitives such as lines and cubes to support schematics. Together, these capabilities streamline the visualization workflow, reducing the time and expertise required to produce publication-ready results.
To encourage adoption and extensibility, we are releasing this library as open source. The Python-based tools are provided as pip-installable packages, allowing researchers to integrate them seamlessly into existing Blender workflows with minimal setup. We demonstrate the library’s use through a series of proof-of-concept examples, ranging from robot motion visualizations to schematic figures and geometric approximations. By providing robotics researchers with accessible scripting interfaces for Blender, our system expands the range of tools available for visual communication and aims to accelerate the integration of high-quality graphics into robotics research.
II Related Works
In this section, we review prior approaches to robotics visualization across three categories: the use of general-purpose game engines, robotics-focused simulators, and specialized visualization tools. By situating our system within this landscape, we highlight the gap between accurate simulation platforms and high-quality rendering tools that our work aims to address.
II-A Game Engines in Robotics
Game engines are widely used in robotics research for visualization and physics simulation [bartneck2015robot, saivichit2024robotic, pukki2021unreal, li2024photorealistic, lixingjian2024photorealistic, wolf2020evolution, young2020unreal, morse2021gaming]. Unity has been employed to support human–robot interaction research through platforms such as the Robot Engine [bartneck2015robot], and to illustrate kinematic concepts in interactive environments [mattingly2012robot]. These works highlight Unity’s strengths in real-time animation, interaction design, and a user-friendly interface.
Unreal Engine has also seen growing adoption, particularly in agricultural robotics and autonomous ground vehicle research. Its ability to simulate messy field conditions, dynamic lighting, and lightweight physics at high frame rates has made it suitable for these domains [li2024photorealistic, lixingjian2024photorealistic, young2020unreal]. However, achieving photorealism typically requires substantial modeling effort and the construction of custom datasets, limiting accessibility for general-purpose robotics visualization.
In contrast to these real-time engines, our system focuses specifically on creating high-quality static renderings and animations. Game engines optimize for interactive performance, often at the expense of lighting fidelity, texture resolution, or ease of setup. By leveraging Blender’s offline rendering capabilities, our library prioritizes photorealism and clarity of presentation rather than simulation speed. This distinction makes it particularly well-suited for generating figures and videos intended for scientific communication.
II-B Simulators in Robotics
Robotic simulators such as Gazebo, CoppeliaSim, Webots, and others are designed to replicate complex robot dynamics and interactions [ivaldi2014tools, de2019analysis, farley2022pick, andreiev2024comparative, nogueira2014comparative]. These tools emphasize accurate modeling of physics, contact, and control, often offering multiple physics engines to balance realism and performance. For example, CoppeliaSim has been shown to accurately capture mobile robot motion across varied terrains by allowing users to select among physics engines [farley2022pick]. Similarly, Gazebo has been favored for humanoid simulation due to its open-source ecosystem, active community, and optimized solvers for multi-body contact [ivaldi2014tools].
While these simulators are indispensable for testing and validation, their visual output is generally secondary to physics fidelity. As a result, rendered images and animations often lack the polish of dedicated graphics tools. Our system complements these simulators by targeting the visualization stage: it imports standard robot descriptions such as URDF, but instead of focusing on accurate physics, it enables researchers to quickly generate clear, visually appealing renderings and animations for communication purposes.
II-C Robot Visualization Tools
Several specialized software packages exist to support robot visualization and education [manseur2004visualization, sadanand2015virtual, gil2015development, miller2005robotic]. For instance, RobotDraw was developed as an interactive platform for teaching manipulator kinematics, allowing students to directly explore virtual robot behavior [manseur2004visualization]. RoboAnalyzer, built on the MATLAB Robotics Toolbox, provides pre-built functions for modeling and analyzing serial-chain robots, lowering the barrier for users without robotics expertise [sadanand2015virtual]. GraspIt! extends these capabilities to robotic grasping, enabling detailed analysis and visualization of hand–object interactions [miller2005robotic].
These systems demonstrate the value of accessible visualization interfaces, but they are typically limited in scope, tied to specific robotics domains, or focused on educational contexts. Visual fidelity is often modest, and their user communities remain relatively small. By embedding robotics-aware functionality into Blender, our system connects roboticists to a large, established computer graphics ecosystem. This integration provides both ease of use and access to advanced rendering features, expanding the possibilities for producing high-quality, publication-ready visualizations.
III Implementation Overview
In this section, we provide an overview of the design and implementation of our system. The goal is to outline the core components, describe how they integrate with Blender, and highlight the scripting interfaces that enable robotics researchers to create high-quality visualizations. We first discuss the overall software structure, then detail the installation and setup process, and finally describe the key scripting functions that form the backbone of the library.
III-A Code Structure and Setup
The library is designed to integrate seamlessly with Blender’s existing Python environment. Installation requires only a single step: the package can be directly pip-installed into the Python interpreter bundled with Blender. No additional dependencies or external toolchains are needed, which keeps the setup lightweight and accessible across different platforms.
Once installed, the workflow is straightforward. Blender already includes a built-in Python scripting environment, so users can directly add short scripts inside Blender or save them as external files to be loaded later. A script can, for example, load a robot specification in URDF format, after which the library automatically parses the URDF, extracts both geometric meshes and kinematic structure, and spawns the robot directly into the Blender scene. From this point, the robot can be configured, animated, or modified using simple scripting functions provided by the library. This organization minimizes boilerplate setup and ensures that researchers can move quickly from installation to generating high-quality visualizations.
A central goal of our library is to make Blender’s rendering and animation capabilities directly accessible to robotics researchers through simple Python scripting. To this end, the library provides a suite of functions organized into three categories: (1) robot and environment importing; (2) configuration and keyframing of robot states and visual attributes; and (3) generation of primitive 3D shapes. These functions are designed to minimize boilerplate code, allowing users to quickly move from importing a robot to producing high-quality visualizations.
III-B Robot Importing
Robots and environments can be imported directly into Blender using standardized URDF descriptions. This process automatically loads both the geometric meshes that define the robot’s appearance and the kinematic structure that governs its configuration. By relying on URDF, researchers can reuse existing specifications without modification, ensuring compatibility with a wide range of commonly used platforms.
A typical script to spawn a robot requires only a few lines of code. The example below demonstrates how a URDF can be loaded and visualized:
IV Evaluation and Results
In this section, we demonstrate the capabilities of our system through a series of examples and benchmarks. The goal is to highlight how the library can be used to generate high-quality visualizations, streamline the process of configuring robots within Blender, and enable dynamic annotations that enrich educational and research figures. In practice, versions of this library have been used to create figures and presentation materials associated with several previous papers[liang2025ad, rakita2018autonomous, rakita2018relaxedik, rakita2018shared, rakita2019remote, rakita2019shared, rakita2019stampede, rakita2020analysis, rakita2020effects, rakita2021collisionik, rakita2021single, rakita2022proxima, rakita2025coherence]. We direct interested readers to these papers to see how our library can be used in real-world academic settings.
IV-A Robot Motion Through Color Gradient
When publishing results, roboticists often need to convey robot motion through a single static image rather than an animation. Our system makes this process straightforward by providing scripting functions that allow motion to be illustrated via smooth color gradients. This approach produces clear, journal-quality figures that effectively communicate dynamic behavior in a compact visual form. Examples are shown in Figures 1 and LABEL:fig:ur5_motion, which highlights how these gradients can be integrated into polished visualizations.
The example in Figure LABEL:fig:ur5_motion was created with fewer than 100 lines of code. After importing the UR5 model, the user only needs to specify the start and end configurations, along with the number of intermediate copies to spawn. The following snippet demonstrates the core logic:
V Discussion
This work aims to lower the barrier to producing clear, publication-quality visuals in robotics by embedding robotics-aware scripting directly within Blender. Rather than competing with simulators or game engines that prioritize physical fidelity or real-time interaction, our library is intentionally designed around three pillars: high-quality offline rendering, a concise Python interface, and compatibility with standardized robot description formats (e.g., URDF).
The examples shown in §IV demonstrate how these design choices translate into practical benefits. Motion-through-color gradients compress temporal behavior into a single image; schematic illustrations overlay lightweight primitives atop robot geometry to communicate mechanical structure or subsystem roles; and approximating geometries such as convex hulls and decompositions can be rendered and styled for pedagogy, collision reasoning, or planning visualization. Because installation targets Blender’s bundled Python using a single pip command, the setup cost is minimal, making it practical for both classroom instruction and collaborative multi-lab projects.
Beyond convenience, a scripting-first workflow significantly improves reproducibility. All figures and videos are generated from code that specifies the exact scene layout, camera configuration, materials, and rendering parameters, enabling version control and seamless regeneration as models or datasets evolve. The same script can produce both stills and animations, ensuring consistency in framing, lighting, and materials across formats. For instructional use, frame-accurate annotations, such as arrows, coordinate axes, and bounding boxes, can be programmatically placed on robot geometry to support step-by-step explanations that emphasize spatial reasoning over software tooling.
The design philosophy emphasizes minimalism and composability. The core API remains small—focusing on import, configuration, keyframing, and primitive generation—so users can build their own extensions (e.g., custom camera paths, lighting presets, annotation utilities) without modifying the library itself. Open-source distribution further supports community contributions, shared templates, and streamlined bug reporting.
V-A Limitations
Our work is subject to specific limitations that highlight opportunities for future development. Primarily, this library functions strictly as a visualization tool, not a physics engine; it does not model dynamics, contact interactions, or closed-loop control. Consequently, interpolation between keyframes is purely kinematic unless driven by external trajectory data. The system also relies on well-formed robot description files, meaning malformed or incomplete URDFs may require manual correction prior to import. regarding performance, high-fidelity scenes, characterized by high-polygon meshes or dense convex decompositions, can increase memory overhead and render times, reflecting the constraints of Blender’s offline renderer. Furthermore, while the library streamlines scene construction, users still require a working knowledge of Blender fundamentals, such as lighting and camera composition, to achieve high-quality results. Finally, installation via pip may vary across operating systems and Blender versions. Ultimately, we design this tool to complement, rather than replace, simulators; it is most effective when paired with external systems for planning and data generation.

### v2

@@
 Our goal is not to replicate the full functionality of physics simulators or game engines, but rather to make it straightforward for researchers to create static images, animations, and schematic illustrations that are both visually appealing and robotics-aware.
-The library centers on three key features: importing robots and environments from standardized descriptions such as URDF, keyframing robot configurations and appearance attributes, and generating common 3D primitives such as lines and cubes to support schematics.
+The library centers on three key features: importing robots and environments from standardized descriptions, keyframing robot configurations and appearance attributes, and generating common 3D primitives such as lines and cubes to support schematics.
 Together, these capabilities streamline the visualization workflow, reducing the time and expertise required to produce publication-ready results.
@@
 III-B Robot Importing
-Robots and environments can be imported directly into Blender using standardized URDF descriptions.
+Robots and environments can be imported directly into Blender using a Universal Robot Description Directory (URDD) [klein2025urdd].
+A URDD can be constructed automatically from a standard URDF.
 This process automatically loads both the geometric meshes that define the robot’s appearance and the kinematic structure that governs its configuration.
-By relying on URDF, researchers can reuse existing specifications without modification, ensuring compatibility with a wide range of commonly used platforms.
-A typical script to spawn a robot requires only a few lines of code. The example below demonstrates how a URDF can be loaded and visualized:
+By relying on URDD, researchers can reuse existing specifications without modification, ensuring compatibility with a wide range of commonly used platforms.
+A typical script to spawn a robot requires only a few lines of code. The example below demonstrates how a URDD can be loaded and visualized:
 IV Evaluation and Results